(differential equations). solve as Bernoulli Equation.

Solve as Bernoulli Ean. y+3y=y
Answer #1

y'+3y=y^4 \\*\\*\implies \frac{dy}{dx}+3y=y^4 \\*\\*\implies \frac1{y^4}\frac{dy}{dx}+3y^{-3}=1-----(1) \\*\\*\text{Putting } y^{-3}=v\implies -3y^{-4}\frac{dy}{dx}=\frac{dv}{dx}\implies \frac1{y^4}\frac{dy}{dx}=-\frac13\frac{dv}{dx} \\*\\*(1) \implies -\frac13\frac{dv}{dx}+3v=1 \\*\\*\implies \frac{dv}{dx}-9v=-3-----(2) \\*\\*\text{Integrating factor=} e^{\int{-9dx}}=e^{-9x} \\*\\*(2) \implies e^{-9x}\frac{dv}{dx}-9ve^{-9x}=-3e^{-9x} \\*\\*\implies \frac{d}{dx}\left(ve^{-9x} \right )=-3e^{-9x} \\*\\*\implies ve^{-9x} =\int{-3e^{-9x}}dx \\*\\*\implies ve^{-9x} =\frac13 e^{-9x} +c, \text{ where c is an arbitrary constant.} \\*\\*\implies v =\frac13 +ce^{9x}=\frac{1+3ce^{9x}}{3} \\*\\*\implies y^{-3} =\frac{1+3ce^{9x}}{3} \\*\\*\implies y^{3} =\frac3{1+3ce^{9x}} \\*\\*\implies y =\sqrt[3]{\frac3{1+3ce^{9x}}}